# BrainECHO: Semantic Brain Signal Decoding through Vector-Quantized Spectrogram Reconstruction for Whisper-Enhanced Text Generation

### v1

@@ Conclusion
This paper introduces a novel three-stage brain-to-text framework, BrainECHO, that addresses the shortcomings of prior methods. These methods relied on teacher forcing and failed to compare model performance against pure noise inputs. BrainECHO bridges the latent spaces of text and corresponding aurally evoked brain signals through vector-quantized spectrogram reconstruction and fine-tuned use of the Whisper model. It achieves state-of-the-art (SOTA) performance on public EEG and MEG datasets across various experimental settings. By extracting deep semantic information from brain signals, BrainECHO provides valuable insights for future research in the brain-to-text decoding paradigm in the BCI field.
Appendix
Appendix A A. Conformer
@@
Appendix B B. Dataset
B.1 Brennan
The Brennan dataset (Brennan and Hale 2019) contains raw electroencephalography (EEG) data collected from 49 human subjects. Participants were asked to passively listen to a 12.4-minute audiobook story of chapter one of Alice’s Advenctures in Wonderland, while their EEG data was recorded. Participants completed an eight-question multiple choice questionnaire concerning the contents of the story at the end of the experimental session. We retain 33 participants’ data who achieved high scores.
The stimulus chapter originally contains 84 sentences. Since the annotation files only provide word-level annotations, directly concatenating words to form sentences would result in the absence of punctuation marks. Therefore, we use WhisperX (Bain et al. 2023) to segment the audio stimulus into segments of no more than 12 seconds, resulting in 140 sentences.
B.2 GWilliams
GWilliams (Gwilliams et al. 2023), known as the “MEG-MASC” dataset, provides raw magnetoencephalography (MEG) data from 27 English speakers who listened to two hours of naturalistic stories. Each participant performed two identical sessions, involving listening to four fictional stories from the Manually Annotated Sub-Corpus (MASC). The four stories are: ‘LW1’ (861 words, 5 min 20 sec), ‘Cable Spool Boy’ (1948 words, 11 min), ‘Easy Money’ (3541 words, 12 min 10 sec) and ‘The Black Willow’ (4652 words, 25 min 50 sec).
To ensure a fair comparison with NeuSpeech (Yang et al. 2024a), we follow its experimental setup by concatenating words with the same sentence ID into full sentences, based on the annotation files. This process results in 661 sentences.
Appendix C C. Dataset Splitting
In this section, we detail the dataset splitting strategies employed in our study. As shown in Table 8, four distinct strategies are utilized, each presenting different levels of evaluation difficulty. The random shuffling strategy is the most basic, incorporating data from all subjects and trials into the training samples. The sentence-based strategy is more challenging, simulating scenarios where samples from different participants are not aligned, resulting in missing data for some sentences for each participant. The session-based and subject-based strategies are the most difficult but also the most realistic, as they assess the model’s ability to generalize to new trials and subjects, respectively. This capability is crucial for the practical application of language-based BCIs. The Brennan dataset utilizes only two splitting methods due to its inclusion of data from a single trial. Consequently, splitting by sentence yields results similar to those obtained by random shuffling.
Appendix D D. Implementation Details
Appendix E E. Evaluation Results
Evaluation metrics of the GWilliams dataset across various splitting strategies are presented in Table 10. NeuSpeech (Yang et al. 2024a), the previous SOTA model for MEG-to-text translation, serves as the baseline for comparison. Given the similarity in data preprocessing and dataset splitting strategies, the NeuSpeech results are cited directly from the original paper. When using random shuffling, BrainECHO achieves a BLEU-4 score of 72.42, outperforming NeuSpeech by 24.64 points (+51.57%). Additionally, with session-based splitting, BrainECHO attains a BLEU-1 score of 75.24, exceeding NeuSpeech by 22.08 points (+41.53%). These results indicate that BrainECHO can generate text that closely matches the ground truth.
Appendix F F. Generated Samples
To intuitively demonstrate the powerful decoding ability of BrainECHO, additional translated examples for the Brennan and GWilliams datasets are presented in Table 11 and 12, respectively.
Appendix G G. Reconstructed Mel Spectrograms
Figure 5 and 6 display some samples of Mel spectrograms reconstructed from the brain signals for the Brennan and GWilliams datasets, respectively. These samples demonstrate that BrainECHO can produces Mel spectrograms that are largely consistent with the ground truth. Notably, the model effectively restores fine details and accurately predicts the intervals and silent segments in the spectrograms. These results highlight the model’s expressive and predictive capabilities, as it can extract Mel spectrograms from brain signal segments exceeding 20 seconds–a feat not achieved by previous methods.
Appendix H H. Codebook Size
To explore the impact of the quantizer, we investigate the performance of BrainECHO with codebook sizes ranging from 1024 to 4096. As shown in Figure 7, the performance peaks at a codebook size of 4096. However, the metrics do not increase linearly with codebook size. When the codebook size increases from 1024 to 2048, the decoding performance improves, but it decreases when the size further increases to 3072. This indicates that a smaller codebook may not capture diverse acoustic representations, while a larger codebook may increase training difficulty and computational burden. Thus, we choose 2048 as the codebook size for balancing performance and efficiency.
@@ B.2 GWilliams
Table 7: The structure of TS Conv. $C$ and $T_{\varepsilon}$ denote the number of EEG/MEG channels and timestamps, respectively.
@@
Brennan Sentence For each participant, 10% of unique sentences are allocated to the test set. The remaining sentences are shuffled and split into train:valid 8:1. Note that the test set for each subject may contain different sentences. 3696:462:462
@@
Configuration P A F P A F
@@
Table 9: Details of the experimental configuration. P, A, F denote the various training stages: pretraining, alignment and finetuning, respectively.
BLEU-N (%) $\uparrow$ ROUGE-1 (%$\uparrow$ WER (%) $\downarrow$
Split Input Method N=1 N=2 N=3 N=4 P R F
Random MEG feature EEG-to-Text (Wang and Ji 2022) 9.21 2.13 0.57 0.14 9.74 10.73 11.38 118.25
Shuffling MEG NeuSpeech (Yang et al. 2024a) 60.3 55.26 51.24 47.78 60.88 59.76 58.73 56.63
MEG BrainECHO 73.35 72.66 72.46 72.42 69.66 70.12 69.73 31.44
Session MEG NeuSpeech (Yang et al. 2024a) 53.16 - - - - - - -
MEG BrainECHO 75.24 74.57 74.34 74.27 72.94 72.84 72.78 29.59
Subject MEG BrainECHO 75.05 74.38 74.18 74.14 71.83 72.02 71.72 29.80
Table 10: Overall comparison of decoding performance on the GWilliams dataset. All methods are evaluated without teacher forcing.
@@
Figure 5: Predicted Mel spectrograms on Brennan dataset.
Figure 6: Predicted Mel spectrograms on the GWilliams dataset. For visualization purposes, only the first half of the Mel spectrograms are displayed.

### v2

@@ 5 Conclusion
This paper introduces a novel three-stage brain-to-text framework, BrainECHO, that addresses the shortcomings of prior methods. These methods relied on teacher forcing and failed to compare model performance against pure noise inputs. BrainECHO bridges the latent spaces of text and corresponding aurally evoked brain signals through vector-quantized spectrogram reconstruction and fine-tuned use of the Whisper model. It achieves SOTA performance on public EEG and MEG datasets across various experimental settings. By extracting deep semantic information from brain signals, BrainECHO provides valuable insights for future research in the brain-to-text decoding paradigm in the BCI field.
Limitations
The limitations of our proposed work are summarized as follows:
Dataset Limitations
While our framework successfully achieves sentence-level decoding constrained by a predefined sentence set, it is still far from higher levels of speech/text decoding — specifically, decoding sentences based on known words or even phonemes from the training set. We conducted preliminary experiments on word-level and phoneme-level generalization using a retrieval-augmented generation approach. However, all of these methods have yielded unsatisfactory decoding performance, with BLEU-1 scores not exceeding 10 and BLEU-4 scores approaching zero. We believe this is largely influenced by the dataset paradigm and the amount of data. Specifically, subjects only passively listen to long continuous stories, lacking engagement in multiple modalities. Additionally, sentence lengths vary significantly, making segmentation challenging, and the vocabulary covered in the corpus is extremely limited and unblanced. These factors hinder the advancement of open-vocabulary decoding paradigms. We encourage future researchers (including ourselves) to collect larger-scale, multimodal datasets with a well-structured stimulus presentation. Such datasets would lay a solid foundation for designing more generalizable and robust decoding frameworks.
Experiment Limitations
In our experimental setting, all data are strictly segmented on a sentence-by-sentence basis before being fed into the model, which may not align with real-world decoding scenarios, due to the potential unknown length of the signals to be translated. Moreover, according to the results reported by NeuSpeech (Yang et al., 2024b), sentence-level decoding may face overfitting issues, as neural signals of different lengths need to be padded to the same length before fed into the model. However, under the condition that there is a correlation between signal length and sentence length, our approach may help the model decode by implicitly injecting the length information of the signal. Moreover, as reported by NeuSpeech (Yang et al., 2024b), sentence-level decoding might encounter overfitting problems. The reason is that neural signals of varying lengths should be padded to a consistent length before being fed into the model. When a correlation exists between signal length and sentence length, it is possible that our proposed approach inadvertently facilitates the model’s decoding by implicitly integrating the length information of the signal. MAD (Yang et al., 2024c) and NeuGPT (Yang et al., 2024a) showed an unsatisfactory result with a uniform signal length, suggesting that the current task of generating open-vocabulary text based solely on the neural signal pattern remains extremely challenging. Our forthcoming research efforts will focus on leveraging LLMs and more efficient alignment strategies to diminish the dependence on length information.
Ethical Statement
This study uses publicly available datasets and does not involve the collection of any brain activity data from human subjects. Therefore, our research does not have any adverse impact on human society.
Acknowledgements
This work is supported by the National Natural Science Foundation of China (Grant No. 62306089), the Shenzhen Science and Technology Program (Grant Nos. RCBS20231211090800003 and ZDSYS20230626091203008) and the Shenzhen-Hong Kong Institute of Brain Science-Shenzhen Fundamental Research Institutions (2023SHIBS0003).
Appendix A Conformer
@@
Table 6: The structure of TS Conv. $C$ and $T_{\varepsilon}$ denote the number of EEG/MEG channels and timestamps.
@@
Brennan Sentence For each participant, sentence-EEG/MEG pairs corresponding to random selected 10% of unique sentences are allocated to the test set, then the remaining sentence-EEG/MEG pairs are shuffled and split into train:valid 8:1. Note that the test set for each subject may contain different sentences and the training set may cover all possible sentences. 3696:462:462
@@
Appendix B Necessity of the Three-Stage Paradigm
Although it is possible to integrate Stage 2 and Stage 3, or even all stages, into a single stage for end-to-end training, we chose to adopt a three-stage training paradigm for several reasons.
First, our three-stage design achieves decoupled representation learning, and each stage serves a distinct and crucial purpose. Stage 1 focuses on optimizing the quantizer, audio encoder and decoder. Stage 2 is dedicated to aligning brain signals with Mel spectrograms. By doing so, we can better capture the complex relationship between the neural activity and the corresponding audio features, which is a key step in bridging the gap between brain signals and text. Finally, in Stage 3, we leverage the Mel spectrograms as an intermediate modality. Through fine-tuning Whisper, we are able to align the brain signals with the text modality. By decoupling brain signal representation learning from linguistic knowledge preservation, we mitigate the risk of LLM over-dominance and establish robust EEG/MEG-to-text mapping.
Second, with only a single Nvidia 3090 (24GB) GPU, we found that it would be challenging to support the large-scale parameter training required for the combined stage. One of the advantages of this three-stage design is that each stage only requires the optimization of a specific subset of module parameters. This significantly reduces the computational burden at each step, enabling us to effectively balance the decoding performance and resource utilization. As demonstrated by our experimental results in Section 4.5, this three-stage training approach is highly effective, highlighting its potential for similar research in the field.
Appendix C Datasets
C.1 Brennan
The Brennan dataset Brennan and Hale (2019) contains raw electroencephalography (EEG) data collected from 49 human subjects. Participants were asked to passively listen to a 12.4-minute audiobook story of chapter one of Alice’s Advenctures in Wonderland, while their EEG data was recorded. Participants completed an eight-question multiple choice questionnaire concerning the contents of the story at the end of the experimental session. We retain 33 participants’ data who achieved high scores.
The stimulus chapter originally contains 84 sentences. Since the annotation files only provide word-level annotations, directly concatenating words to form sentences would result in the absence of punctuation marks. Therefore, we use WhisperX Bain et al. (2023) to segment the audio stimulus into segments of no more than 12 seconds, resulting in 140 sentences.
C.2 GWilliams
GWilliams Gwilliams et al. (2023), known as the “MEG-MASC” dataset, provides raw magnetoencephalography (MEG) data from 27 English speakers who listened to two hours of naturalistic stories. Each participant performed two identical sessions, involving listening to four fictional stories from the Manually Annotated Sub-Corpus (MASC). The four stories are: ‘LW1’ (861 words, 5 min 20 sec), ‘Cable Spool Boy’ (1948 words, 11 min), ‘Easy Money’ (3541 words, 12 min 10 sec) and ‘The Black Willow’ (4652 words, 25 min 50 sec).
To ensure a fair comparison with NeuSpeech Yang et al. (2024b), we follow its experimental setup by concatenating words with the same sentence ID into full sentences, based on the annotation files. This process results in 661 sentences.
@@ C.2 GWilliams
Configuration Pretraining Alignment Finetuning Pretraining Alignment Finetuning
@@
Table 8: Details of the experimental configuration.
Figure 5: Predicted Mel spectrograms on the GWilliams dataset.
Generated samples on Brennan
EEG-to-Text But they were all locked, and when Alice had been all the way down one side and up the other trying every door, she did not care how she was ever to get out again.
NeuSpeech There seemed to be no use in waiting by the little door, so she went back to the table.
EEG-to-Text how she longed to get out of that dark hall and wander about among those beds of bright flowers and those cool fountains, but she did not even get her head through the doorway.
NeuSpeech But they were all locked, and when Alice had been all the way down one side and up the other trying every door, she walked sadly down the middle, wondering how she was ever to get out again.
Generated samples on GWilliams
EEG-to-Text It was a long time since I had last seen him in the flesh
NeuSpeech I seen him since high school when I was young, at least before and we were never buddies in any place.
EEG-to-Text He said he had no idea how long it would take him to get back home
NeuSpeech My patience was long gone and I was back in the car. But when I heard that many of you were looking for whatever it was, but what about this?
Table 9: Comparison of decoding sentences generated by different methods, where bold and underline indicate an exact match and a similar match, respectively, between prediction and ground truth. All methods use the same generation configuration. w/o ft means decoding by inputting the predicted Mel spectrogram into Whisper directly without fine-tuning in the final stage. Only examples of NeuSpeech are reported rather than those of MAD because of NeuSpeech’s overall superior performance and the similarity of its method to MAD’s.
@@
Appendix D Dataset Splitting
In this section, we detail the dataset-splitting strategies employed in our study. As shown in Table 7, four distinct strategies are utilized, each presenting different levels of evaluation difficulty. The random shuffling strategy is the most basic, incorporating data from all subjects and trials into the training samples. The sentence-based strategy is more challenging, simulating scenarios where samples from different participants are not aligned, resulting in missing data for some sentences for each participant. The session-based and subject-based strategies are the most difficult but also the most realistic, as they assess the model’s ability to generalize to new trials and subjects, respectively. This capability is crucial for the practical application of language-based BCIs. The Brennan dataset utilizes only two splitting methods due to its inclusion of data from a single trial. Consequently, splitting by sentence yields results similar to those obtained by random shuffling.
Appendix E Implementation Details
Appendix F Examples of Generated Sentences
A selection of samples generated from different methods are shown in Table 9. These examples indicate that BrainECHO can produce sentences that closely match the original text, even when the reference is long and intricate. Remarkably, even without the final fine-tuning of Whisper, BrainECHO still generates results highly relevant to the original text, highlighting the effectiveness of brain-audio latent space alignment (stage 2). In contrast, EEG-to-Text (Wang and Ji, 2022) experiences difficulties in generating semantically relevant sentences, and NeuSpeech (Yang et al., 2024b) may generate content unrelated to the ground truth when decoding long sentences, which can have a significant impact on practical applications in high-precision decoding scenarios.
To intuitively demonstrate the powerful decoding ability of BrainECHO, additional translated examples for the Brennan and GWilliams datasets are presented in Table 10 and 11, respectively. For most test samples, our method demonstrates accurate decoding. However, for certain samples, our model generates completely unrelated content, such as "There were doors all around the hall." and "What a curious feeling, said Alice." in Table 10. This suggests that the model may struggle with discriminability in sentences of similar length, highlighting the persistent challenge of extracting semantically relevant patterns from low signal-to-noise non-invasive signals.
Some samples of Mel spectrograms reconstructed from the brain signals for the GWilliams datasets are shown in Figure 5.
@@ Appendix F Examples of Generated Sentences
Appendix G Additional Experiments on Codebook Size in the Quantizer
To further explore the impact of the quantizer, we investigate the performance of BrainECHO with codebook sizes ranging from 1024 to 4096. As shown in Figure 6, the performance peaks at a codebook size of 4096. However, the metrics do not increase linearly with codebook size. When the codebook size increases from 1024 to 2048, the decoding performance improves, but it decreases when the size further increases to 3072. This indicates that a smaller codebook may not capture diverse acoustic representations, while a larger codebook may increase training difficulty and computational burden. Thus, we choose 2048 as the codebook size for balancing performance and efficiency.
Appendix H The Reason Why the Test Results Make Sense
Since we split the data in paired form (i.e., "Semantic Audio – Brain Signal evoked by the Semantic Audio"), there could be cases where the same sentence, but with different brain signals from different people (e.g., Sentence $t$ – Brain Signal $\varepsilon_{subj1}$, Sentence $t$ – Brain Signal $\varepsilon_{subj2}$), is included in the training or test set. Therefore, the Mel spectrograms during stage 1 could have already been seen during training, even though they are part of the test set.
However, for non-invasive natural language brain-computer interfaces, unlike invasive systems that decode neural activity related to language-specific motor areas (Willett et al., 2023; Metzger et al., 2023; Chen et al., 2024b), non-invasive interfaces have lower signal-to-noise ratios. Furthermore, similar to invasive systems, decoding of brain signals occurs on previously seen sentences, with the vocabulary expanding progressively to achieve open vocabulary. Testing with completely unseen sentences can be overly ambitious, as demonstrated in NeuSpeech Yang et al. (2024b), where testing on completely unseen sentences resulted in a BLEU-1 score of only 6.91. Like the baseline in the paper, we ensure that the same sentence and its evoked brain signal do not appear in both the training and testing stages. Additionally, we have tested various data split scenarios (session, sentence and subject).
Moreover, to prevent data leakage, even if Mel spectrograms from the test set were exposed during stage 1, the brain signals from the test set were never used in stage 2. Stage 1 only serves to obtain a low-dimensional representation of the Mel spectrograms, akin to creating a feature selector for Mel spectrograms. The brain signals decoded during testing are always from data that was not seen during training.